Fill in the P(X=x)

values in the table below to give a legitimate probability distribution for the discrete random variable X, whose possible values are

1,3,4,5, and 6

Value x of X

P (X=x)

1

3

4

0.28

5

0.15

6

0.23
Answer #1

Answer )

We know that sum of all the probabilities is equal to 1

Let the probability for 1 and 3 be x and y respectively

X + y + 0.28 + 0.15 + 0.23 = 1

X + Y = 1 - (0.28 + 0.15 + 0.23)

X + Y = 0.34

So, two possible values are 0.17 + 0.17 = 0.34

X = 0.17

Y = 0.17